# A possible statistics loophole in Bell’s theorem

### v2

@@ Abstract
Bell’s theorem proves the incompatibility between quantum mechanics and local realistic hidden-variable theories. In this paper we first show that, contrary to a common belief, the theoretical proof of Bell’s theorem is not affected by counterfactual reasoning. Second, we demonstrate that the experimental verification of this theorem may be affected by our unavoidable ignorance about the probability distribution of the hidden variables. Our study is based on the standard theory of random variables, and lays the groundwork for a critical rethinking of Bell’s theorem and its consequences.
@@ I Introduction
Quantum mechanics is at the same time the most successful and the most debated theory in physics we have, still today more than one hundreds years after its foundation. One of the reasons for this debate originated from the celebrated paper by Einstein, Podolsky and Rosen [1], where the authors discussed the possible incompleteness of quantum mechanics (QM), and the need to complete it by introducing additional variables, which later became known as hidden variables. The rest is well-known history and needs not be repeated here. We only point out that the turning point in that history was represented by the so-called Bell’s theorem and its by-products [2]. In short, Bell proved that no local, realistic hidden-variable (HV) theory could describe the outcomes of certain correlation experiments with entangled particle [3, 4]. He was able to express his theorem in the form of an inequality that could be verified experimentally. Since then, many variants of such inequality have been derived [5], like the largely used CHSH (Clauser, Horne, Shimony et Holt) inequality [6]. More recently, Bell’s theorem has made a comeback mainly due to its usage in the field of quantum information [7]. Despite these successes, Bell’s theorem has not been free from criticisms [8, 9, 10, 11, 12, 13], the most significant of which is perhaps the use of counterfactual reasoning [14].
The purpose of this work is twofold. First, we demonstrate that, in fact, there is no counterfactual reasoning in the proof of the CHSH inequality. Then, we show that any experimental verification of this inequality is inevitably affected by a loophole rooted in our ignorance about the statistical distribution of the hidden variables. The tools we use to obtain our results are simply basic quantum mechanics and random variable theory.
This paper is structured as follows. In Sec. II, we present a quick review of two-photon correlation experiments. Next, in Sec. III we show that counterfactual reasoning is not necessary to prove the CHSH inequality. This inequality is usually written in terms of the so-called Bell parameter $S$, as $\left|S\right|\leq 2$. The parameter $S$ is given by the mean value of a suitably defined random variable $T$, which describes a specific linear combination of the outcomes of four distinct correlation experiments. In Sec. IV we perform a detailed statistical analysis of $T$ showing that the sample mean of $T$, which is the quantity actually measured in real-world experiments, may differ from $S$ in an unknowable way. Then, in Sec. V we calculate the actual upper bound of CHSH’s inequality. Finally, in Sec. VI, we draw our conclusions. Four appendixes reporting detailed calculations, complete the paper.
@@ II A brief review of two-photon correlation experiments
where the two vectors $|e_{k}^{(A)}\rangle,~{}(k=\pm 1)$, form a complete and orthonormal basis in the two-dimensional Hilbert space $\mathcal{H}_{A}$ representing the polarization state of photon $A$. The same definitions hold, mutatis mutandis, for photon $B$. By definition, $|\psi\rangle\in\mathcal{H}=\mathcal{H}^{(A)}\otimes\mathcal{H}^{(B)}$, where $\mathcal{H}$ is the four-dimensional Hilbert space of the pair of photons. Note that throughout this paper, when an index takes on the two values ±$\pm 1$, we write $q_{\pm}$ as an abbreviation of $q_{\pm 1}$, for the generic quantity $q_{k},~{}(k=\pm 1)$.
@@
where $\alpha\in[0,\pi)$ is the angle by which the analyzer is rotated. Conversely, when the photon exits from the port $d_{A-}$, its polarization is represented by the vector state
@@
Again, the same definitions apply to photon $B$ replacing everywhere $A$ with $B$ and $\alpha$ with $\beta$. Behind each of Alice and Bob’s output ports of their analyzers there is a detector that fires when reached by a photon.
In this context, the term “experiment” denotes the emission of a pair of photons by the light source, followed by the firing of a pair of Alice’s and Bob’s detectors. The possible outcomes of this experiment are given by the set $\Omega$ (sample space of the experiment), whose elements are all the ordered pairs $(d_{Ak},d_{Bl}){\,\vcentcolon=\,}\omega_{kl},~{}(k,l=\pm 1)$, that is
@@
where $+$ and -- stand for $+1$ and $-1$, respectively. Then, for example, if we record the outcome $(d_{A+},d_{B-})$, this means that Alice’s detector behind port $d_{A+}$ has fired, and Bob’s detector behind $d_{B-}$, has also fired. We remark that each experiment, which we denote by $E$, is characterized by an ordered pair of angles $(\alpha,\beta)$ that describe the orientation of Alice and Bob’s analyzers respectively, hence we write $E=E(\alpha,\beta)$. Different pairs of angles characterize different experiments, so that $E(\alpha,\beta)\neq E(\alpha^{\prime},\beta^{\prime})$, when $(\alpha,\beta)\neq(\alpha^{\prime},\beta^{\prime})$.
@@
with $k,l=\pm 1$, and here and hereafter $\mathbb{E}[U]$ denotes the expected value of the random variable $U$.
@@
This expression is valid irrespective of whatever theory we use to calculate $C(\alpha,\beta)$, either quantum mechanics or a hidden-variable theory. The only obvious constraint to $C(\alpha,\beta)$ is
@@
The four probabilities $p_{kl}$ can be estimated from the experimental data by repeating the experiment $N\gg 1$ times and using the frequency interpretation of probability [16]. Let $N_{kl},~{}(k,l=\pm 1)$ be the number of simultaneous clicks of the detectors behinds port $d_{Ak}$ and $d_{Bl}$. Then we estimate
@@
Alternatively, the correlation coefficient $\mathbb{E}[XY]$ can be calculated by using either a) quantum mechanics, or b) some hypothetical hidden-variable theories. Bell’s theorem [5] basically consists of comparing certain linear combinations of correlation coefficients calculated according to a) and b). A popular form of Bell’s theorem is the so-called CHSH’s inequality. It is often stated in textbooks that counterfactual reasoning is required to derive this inequality [14, 17]. In the next section we will show that this belief is unjustified.
@@ III.1 General facts
The probability distribution $p_{kl}(\alpha_{\mu},\beta_{\mu})$ of the outcomes of the experiment $E_{\mu}(\alpha_{\mu},\beta_{\mu})$ can be calculated either using quantum mechanics, thus obtaining $p_{kl}^{\text{QM}}(\alpha_{\mu},\beta_{\mu})$, or local realistic hidden-variable theories which supposedly give $p_{kl}^{\text{HV}}(\alpha_{\mu},\beta_{\mu})$. What is important to remark here is that in both cases, since the four experiments $E_{1},\ldots,E_{4}$ are performed independently, the joint probability distribution of the eight random variables $X_{1},Y_{1},X_{2},Y_{2},X_{3},Y_{3},X_{4},Y_{4}$, is given by the product of the individual distributions of the four experiments, viz.,
@@
In Eq. (III.1) $f_{\Lambda}(\lambda){\,\vcentcolon=\,}\rho(\lambda)$ denotes the probability density function of the random vector $\Lambda$, characterized by the domain $\mathcal{S}$ and the range $\mathcal{D}$:
@@
for all $\lambda\in\mathcal{D}$. In the remainder we will use both notations according to convenience. These functions are what mathematicians call simple functions [18], that is measurable functions that takes only finitely many values. They can be written as,
@@ III.2 The CHSH inequality
From this equation it is clear that the quantity $ab+ab^{\prime}+a^{\prime}b-a^{\prime}b^{\prime}$ does not need to represent a physically measurable quantity, it is just the result of a formal mathematical manipulation that we performed to pass from (40) to (41). However, in the traditional derivations of the CHSH inequality presented in many textbooks and articles, such linearity is used in reversed manner, that is they first calculate $M(\lambda)=a(\lambda)b(\lambda)+a(\lambda)b^{\prime}(\lambda)+a^{\prime}(\lambda)b(\lambda)-a^{\prime}(\lambda)b^{\prime}(\lambda)$, which is physically not measurable, and then evaluate the average
@@ IV The Bell parameter as a random variable
We can find the probability density function $f_{T}(t)$ of $T$ using the random variable transformation theorem [22]. A straightforward calculation gives
@@
with $\delta(i,j)$ denoting the Kronecker delta function
@@
$\phantom{=4.}-C^{2}(\alpha^{\prime},\beta)-C^{2}(\alpha^{\prime},\beta^{\prime}),$ (55)
@@
Now, suppose to run the series of four experiments $\{E_{\mu}\}=\{E_{1},\,E_{2},\,E_{3},\,E_{4}\}$ a number $N$ of times. In this case we can estimate the mean $\mathbb{E}\bigl{[}T\bigr{]}$ and the variance $\mathbb{V}\bigl{[}T\bigr{]}$ of $T$, by sampling $T$ $N$ times instead of performing the integrations or the sums in (IV) and (IV). Let $t(n)$ be the value of $T$ obtained in the $n\text{th}$ run of the series of four experiments $\{E_{\mu}\}$. Then, the sample mean
@@
Thus, $\mu_{N}$ is simply the $N$-sample estimation of the Bell parameter $S=\mu=\mathbb{E}[T]$ defined by (III.1). Equation (45) tells us that $\left|S\right|\leq 2$ according to HV theories, and from the law of large numbers [24], we know that $\mu_{N}\to\mu$ for $N$ sufficiently large. Therefore, in such a limit, we expect that, approximately, $\left|\mu_{N}\right|\leq 2$. However, since all real-world experiments can be repeated only a finite number $N$ of times, the key question is: how large must be $N$ to guarantee the validity of $\left|\mu_{N}\right|\leq 2$? In other words, how likely is it to find $\left|\mu_{N}\right|>2$ for any finite $N$, in the case that the physics of the four experiments $\{E_{\mu}\}$ is correctly described by a hidden variable theory? In the remainder we are going to show that it is impossible to answer this question, due to our ignorance about the random hidden variables $\lambda$, and their distribution $\rho(\lambda)$.
@@ IV.1 Calculating the probability distribution of $\mu_{N}$
From this definition it follows that $\mu_{N}$ is one of the values taken by $M_{N}$ in a run of the series of $4N$ experiments. Using again the random variable transformation theorem [22], we can calculate the probability density function $f_{M_{N}}(m)$ of $M_{N}$. A straightforward calculation gives
@@
where $t_{i_{n}}$ is defined by (50). From this equation it follows that the probability to find, after a series of $4N$ experiments, $\mu_{N}$ bigger than a certain value, say $m_{0}>0$, is given by
@@
where $\text{N}(\mu,\sigma^{2}/N)$ denotes the normal distribution with mean $\mu$ and variance $\sigma^{2}/N$, and the latter two parameters are calculated using (IV) and (IV), respectively. In this limit of large $N$, the probability to find, after a series of $4N$ experiments, $\mu_{N}\geq m_{0}>0$, is approximately given by
@@ IV.2 A simple example
Next, we choose the a set of angles $\{\alpha,\beta,\alpha^{\prime},\beta^{\prime}\}$, such that the violation of CHSH’s inequality is maximal,
@@
Let us choose $N=10$ and calculate the probability that $\mu_{N}^{\text{QM}}\geq 3>\mu^{\text{QM}}$ using both the exact formula (IV.1), and the approximate one (63). A straightforward calculation gives
@@
We observe that in this example the QM and HV theories were treated equally, in the sense that in both cases we knew the probability distributions of both $\mu_{N}^{\text{QM}}$ and $\mu_{N}^{\text{HV}}$. However, this is not the case in the actual experiments, where we ignore the details of the HV theories. In essence, it is in this disparity of statistical knowledge between quantum and hidden variables that is rooted the problem of verifying Bell’s theorem studied in this work.
@@ V Actual bounds on the sample mean of the Bell parameter $S$
In the previous section we have seen that HV theories permit to calculate the mean value $S$ of the random variable $T$, but not the sample mean $\mu_{N}$. Hidden variables theories also furnish the bounds $-2\leq S\leq 2$. Therefore, it is natural to wonder what we can say about the bounds of $\mu_{N}$. That is what we will do in this section.
@@
$\{\lambda_{\mu,1},\lambda_{\mu,2},\ldots,\lambda_{\mu,N}\},$ (82)
@@
The difference between $M_{n}$ in (V) and $M(\lambda)$ in (41), is that while $M(\lambda)$ depends on a single set of values $\lambda$, $M_{n}$ depends on the four set of values $\lambda_{1,n},\lambda_{2,n},\lambda_{3,n}$ and $\lambda_{4,n}$. Since the latter are independent numbers sampled from the same probability distribution $\rho(\lambda)$, we have that

### v3

@@ Abstract
Bell’s theorem proves the incompatibility between quantum mechanics and local realistic hidden-variable theories. In this paper we show that, contrary to a common belief, the theoretical proof of Bell’s theorem is not affected by counterfactual reasoning. Then, we demonstrate that the experimental verification of this theorem may be affected in an unknowable way by our ignorance about the probability distribution of the hidden variables. Our study is based on the standard theory of random variables, and lays the groundwork for a critical rethinking of Bell’s theorem and its consequences.
@@ I Introduction
Quantum mechanics is at the same time the most successful and the most debated theory in physics we have, still today more than one hundreds years after its foundation. One of the reasons for this debate originated from the celebrated paper by Einstein, Podolsky and Rosen [1], where the authors discussed the possible incompleteness of quantum mechanics (QM), and the need to complete it by introducing additional variables, which later became known as hidden variables. The rest is well-known history and needs not be repeated here. We only point out that the turning point in that history was represented by the so-called Bell’s theorem and its by-products [2]. In short, Bell proved that no local, realistic hidden-variable (HV) theory could describe the outcomes of certain correlation experiments with entangled particle [3, 4]. He was able to express his theorem in the form of an inequality that could be verified experimentally. Since then, many variants of such inequality have been derived [5], like the largely used Clauser, Horne, Shimony and Holt (CHSH) inequality [6]. More recently, Bell’s theorem has made a comeback mainly because of its usage in the field of quantum information [7]. Despite these successes, Bell’s theorem has not been free from criticisms [8, 9, 10, 11, 12, 13], the most significant of which is perhaps the use of counterfactual reasoning [14, 15].
The purpose of this work is twofold. First, we demonstrate that, in fact, there is no counterfactual reasoning in the proof of CHSH’s inequality. Then, we show that any experimental verification of this inequality is inevitably affected by a loophole rooted in our ignorance about the statistical distribution of the hidden variables. The tools we use to obtain our results are simply basic quantum mechanics and random variable theory.
This paper is structured as follows. In Sec. II, we present a quick review of two-photon correlation experiments. Next, in Sec. III we show that counterfactual reasoning is not necessary to prove the CHSH inequality. This inequality is usually written in terms of the so-called Bell parameter $S$, as $\left|S\right|\leq 2$. The parameter $S$ is given by the mean value of a suitably defined random variable $T$, which describes a specific linear combination of the outcomes of four distinct correlation experiments. In Sec. IV we perform a detailed statistical analysis of $T$ showing that the sample mean of $T$, which is the quantity actually measured in real-world experiments, may differ from $S$, the mean value of $T$, in an unknowable way. Then, in Sec. V we calculate the actual upper bound of CHSH’s inequality. Finally, in Sec. VI, we draw our conclusions. Four appendixes reporting detailed calculations, complete the paper.
@@ II A brief review of two-photon correlation experiments
where the two vectors $|e_{k}^{(A)}\rangle,~{}(k=\pm 1)$, form a complete and orthonormal basis in the two-dimensional Hilbert space $\mathcal{H}_{A}$ representing the polarization state of photon $A$. The same definitions hold, mutatis mutandis, for photon $B$. By construction, $|\psi\rangle\in\mathcal{H}=\mathcal{H}^{(A)}\otimes\mathcal{H}^{(B)}$, where $\mathcal{H}$ is the four-dimensional Hilbert space of the pair of photons. Note that throughout this paper, when an index takes on the two values ±$\pm 1$, we write $q_{\pm}$ as an abbreviation of $q_{\pm 1}$, for the generic quantity $q_{k},~{}(k=\pm 1)$.
@@
where $\alpha\in[0,\pi)$ is the angle by which Alice’s polarization analyzer is rotated. Conversely, when the photon exits from the port $d_{A-}$, its polarization is represented by the vector state
@@
Again, the same definitions apply to photon $B$ replacing everywhere $A$ with $B$ and $\alpha$ with $\beta$. Behind each of Alice and Bob’s output ports of their analyzers, there is a detector that fires when reached by a photon.
In this context, the term “experiment” denotes the emission of a pair of photons by the light source, followed by the firing of a pair of Alice’s and Bob’s detectors. The possible outcomes of this experiment are given by the set $\Omega$ (sample space of the experiment), whose elements are all the ordered pairs $(d_{Ak},d_{Bl}){\,\vcentcolon=\,}\omega_{kl},~{}(k,l=\pm 1)$, defined by
@@
where $+$ and -- stand for $+1$ and $-1$, respectively. Then, for example, the outcome $\omega_{+-}=(d_{A+},d_{B-})$ describes the event where Alice’s detector behind port $d_{A+}$ has fired, and Bob’s detector behind $d_{B-}$, has also fired. We remark that each experiment, which we denote by $E$, is characterized by an ordered pair of angles $(\alpha,\beta)$, which describes the orientation of Alice and Bob’s analyzers respectively; hence we write $E=E(\alpha,\beta)$. Different pairs of angles characterize different experiments, so that $E(\alpha,\beta)\neq E(\alpha^{\prime},\beta^{\prime})$, when $(\alpha,\beta)\neq(\alpha^{\prime},\beta^{\prime})$.
@@
with $k,l=\pm 1$, and here and hereafter $\mathbb{E}[U]$ denotes the expectation value of the random variable $U$.
@@
This expression is valid irrespective of whatever theory we use to calculate $C(\alpha,\beta)$, either quantum mechanics or a hidden-variable theory. The only constraint to $C(\alpha,\beta)$ is the normalization
@@
The four probabilities $p_{kl}$ can be estimated from the experimental data by repeating the experiment $N\gg 1$ times and using the frequency interpretation of probability [17]. Let $N_{kl},~{}(k,l=\pm 1)$ be the number of simultaneous clicks of the detectors behinds port $d_{Ak}$ and $d_{Bl}$, after $N$ runs of the experiment. Then, we estimate
@@
Alternatively, the correlation coefficient $\mathbb{E}[XY]$ can be calculated by using either a) quantum mechanics, or b) some hypothetical hidden-variable theory. Bell’s theorem [5] basically consists of comparing certain linear combinations of correlation coefficients calculated according to a) and b). A popular form of Bell’s theorem is the so-called CHSH’s inequality. It is often stated in textbooks that counterfactual reasoning is required to derive this inequality [14, 18]. In the next section we will show that this belief is unjustified.
@@ III.1 General facts
The probability distribution $p_{kl}(\alpha_{\mu},\beta_{\mu})$ of the outcomes of the experiment $E_{\mu}(\alpha_{\mu},\beta_{\mu})$ can be calculated either using quantum mechanics, thus obtaining $p_{kl}^{\text{QM}}(\alpha_{\mu},\beta_{\mu})$, or local realistic hidden-variable theories which supposedly give $p_{kl}^{\text{HV}}(\alpha_{\mu},\beta_{\mu})$. What is important to remark here is that in both cases, since the four experiments $E_{1},E_{2},E_{3}$, and $E_{4}$ are supposedly performed independently, the joint probability distribution of the eight random variables $X_{1},Y_{1},X_{2},Y_{2},X_{3},Y_{3},X_{4},Y_{4}$, is given by the product of the individual distributions of the outcomes of the four experiments, viz.,
@@
In Eq. (III.1) $f_{\Lambda}(\lambda){\,\vcentcolon=\,}\rho(\lambda)$ denotes the probability density function of the random vector $\Lambda$ describing the hidden variables, which has domain $\mathcal{S}$ and range $\mathcal{D}$:
@@
for all $\lambda\in\mathcal{D}$. These functions are what mathematicians call simple functions [19], that is measurable functions that takes only finitely many values. They can be written as,
@@ III.2 The CHSH inequality
From this equation it is clear that the quantity $ab+ab^{\prime}+a^{\prime}b-a^{\prime}b^{\prime}$ does not need to represent a physically measurable quantity, because is just the result of the formal mathematical manipulation that we performed to pass from (40) to (41). However, in the traditional derivations of the CHSH inequality presented in many textbooks and articles, such linearity is used in reversed manner, that is they first calculate $M(\lambda)=a(\lambda)b(\lambda)+a(\lambda)b^{\prime}(\lambda)+a^{\prime}(\lambda)b(\lambda)-a^{\prime}(\lambda)b^{\prime}(\lambda)$, which is physically not measurable, and then evaluate the average
@@ IV The Bell parameter as a random variable
We can find the probability density function $f_{T}(t)$ of $T$, using the random variable transformation theorem [23]. A straightforward calculation gives
@@
with $i=1,\ldots,5$, and $\delta(i,j)$ denotes the Kronecker delta function
@@
$\phantom{=.}-C^{2}(\alpha^{\prime},\beta)-C^{2}(\alpha^{\prime},\beta^{\prime}),$ (55)
@@
Now, suppose to run the series of four experiments $\{E_{\mu}\}=\{E_{1},\,E_{2},\,E_{3},\,E_{4}\}$ $N$ of times. In this case we can estimate the mean $\mathbb{E}\bigl{[}T\bigr{]}$ and the variance $\mathbb{V}\bigl{[}T\bigr{]}$ of $T$, by sampling $T$ $N$ times instead of performing the integrations or the sums in (IV) and (IV). Let $t(n)$ be the value of $T$ obtained in the $n\text{th}$ run of the series of four experiments $\{E_{\mu}\}$. Then, the sample mean
@@
Thus, $\mu_{N}$ is simply the $N$-sample estimation of the Bell parameter $S=\mu=\mathbb{E}[T]$ defined by (III.1). Equation (45) tells us that $\left|S\right|\leq 2$ according to HV theories, and from the law of large numbers [25], we know that $\mu_{N}\to\mu$ for $N$ sufficiently large. Therefore, in such a limit, we expect that, approximately, $\left|\mu_{N}\right|\leq 2$. However, since all real-world experiments can be repeated only a finite number $N$ of times, the key question is: how large must be $N$ to guarantee the validity of $\left|\mu_{N}\right|\leq 2$? In other words, how likely would be to find $\left|\mu_{N}\right|>2$ for any finite $N$, in the case that the physics of the four experiments $\{E_{\mu}\}$ were correctly described by a hidden-variable theory? In the remainder we are going to show that it is impossible to answer this question, due to our ignorance about the random hidden variables $\lambda$, and their distribution $\rho(\lambda)$.
@@ IV.1 Calculating the probability distribution of $\mu_{N}$
From this definition it follows that $\mu_{N}$ is a value taken by $M_{N}$ in a run of the series of four experiments $\{E_{\mu}\}$. Using again the random variable transformation theorem [23], we can calculate the probability density function $f_{M_{N}}(m)$ of $M_{N}$. A straightforward calculation gives
@@
where $t_{i_{n}}$ is defined by (50). From this equation it follows that the probability to find, after performing $N$ times the series of four experiments $\{E_{\mu}\}$, $\mu_{N}$ bigger than a certain value, say $m_{0}>0$, is given by
@@
where $\text{N}(\mu,\sigma^{2}/N)$ denotes the normal distribution with mean $\mu$ and variance $\sigma^{2}/N$, and the latter two parameters are calculated using (IV) and (IV), respectively. In this limit of large $N$, the probability to find $\mu_{N}\geq m_{0}>0$, is approximately given by
@@ IV.2 A simple example
Next, we choose the set of angles $\{\alpha,\beta,\alpha^{\prime},\beta^{\prime}\}$, such that the violation of CHSH’s inequality is maximal,
@@
Let us choose $N=10$, and calculate the probability that $\mu_{N}^{\text{QM}}\geq 3>\mu^{\text{QM}}$ using both the exact formula (IV.1), and the approximate one (63). A straightforward calculation gives
@@
We remark that in this example the QM and HV theories were treated equally, in the sense that in both cases we knew the probability distributions of both $\mu_{N}^{\text{QM}}$ and $\mu_{N}^{\text{HV}}$. However, this is not the case in the actual experiments, where we ignore the details of the HV theories. In essence, it is in this disparity of statistical knowledge between quantum and hidden variables that is rooted the problem of verifying Bell’s theorem studied in this work.
@@ V Actual bounds on the sample mean of the Bell parameter $S$
In the previous section we have seen that HV theories permit to calculate the mean value $S=\mu$ of the random variable $T$, but not the sample mean $\mu_{N}$. Hidden variables theories also furnish the bounds $-2\leq S\leq 2$. Therefore, it is natural to wonder what we can say about the bounds of $\mu_{N}$. That is what we will do in this section.
@@
$\Lambda_{\mu}=\{\lambda_{\mu,1},\lambda_{\mu,2},\ldots,\lambda_{\mu,N}\},$ (82)
@@
The difference between $M_{n}$ in (V) and $M(\lambda)$ in (III.2), is that while $M(\lambda)$ depends on a single set of values $\lambda$, $M_{n}$ depends on the four set of values $\lambda_{1,n},\lambda_{2,n},\lambda_{3,n}$ and $\lambda_{4,n}$. Since the latter are independent numbers sampled from the same probability distribution $\rho(\lambda)$, we have that